# On the Spectral Clustering of a Class of Multigrid Preconditioners

### v2

@@ 3.2 Invariant subspaces – the general case
In this section, we show that the ideas of the simplified case in Section 3.1 extend to more general setting. We assume only that $n\geq N/2$ and $T$ is diagonalizable. Zero may be an eigenvalue of $T$, and we denote its multiplicity by $k$. Eigenvectors with eigenvalue zero are denoted $\mathbf{w}_{\lambda=0}$, while all others are denoted $\mathbf{w}_{\lambda\neq 0}$.
@@
This implies that $\mathtt{E}_{\lambda\neq 0}$ is also rank-one and only one of its two eigenvalues is nonzero. Apply $\mathtt{E}_{\lambda\neq 0}$ to the vector $\mathtt{E}^{(s,\mathrm{full})}_{\lambda\neq 0}\mathbf{u}$:
(14) $\mathtt{E}_{\lambda\neq 0}\mathtt{E}^{(s,\mathrm{full})}_{\lambda\neq 0}\mathbf{u}$ $=\left(\mathtt{E}^{(s,\mathrm{full})}_{\lambda\neq 0}\mathbf{u}\mathbf{v}(\lambda)^{*}\mathtt{E}^{(s,\mathrm{full})}_{\lambda\neq 0}\right)\mathtt{E}^{(s,\mathrm{full})}_{\lambda\neq 0}\mathbf{u},$
@@
For each eigenpair with zero eigenvalue $(\lambda,\mathbf{w}_{\lambda\neq 0})$ of $T$ define the two-dimensional subspace $\mathcal{V}_{\lambda\neq 0}\subset\mathbb{C}^{N}$ through
(16) $V_{\lambda=0}:=\begin{bmatrix}\mathbf{0}&\mathbf{w}_{\lambda=0}\\ D^{-1}C\mathbf{w}_{\lambda=0}&\mathbf{0}\end{bmatrix}\quad\mathrm{and}\quad\mathcal{V}_{\lambda=0}=\mathrm{Im}\left(V_{\lambda=0}\right).$
@@
$E^{(s)}V_{\lambda=0}$ $=\begin{bmatrix}(1-\alpha)I&-\alpha A^{-1}B\\ -\alpha D^{-1}C&(1-\alpha)I\end{bmatrix}\begin{bmatrix}\mathbf{0}&\mathbf{w}_{\lambda=0}\\ D^{-1}C\mathbf{w}_{\lambda=0}&\mathbf{0}\end{bmatrix}$
@@
$E^{(c)}V_{\lambda=0}$ $=\begin{bmatrix}I&A^{-1}B\\ &\end{bmatrix}\begin{bmatrix}\mathbf{0}&\mathbf{w}_{\lambda=0}\\ D^{-1}C\mathbf{w}_{\lambda=0}&\mathbf{0}\end{bmatrix}=V_{\lambda=0}\begin{bmatrix}0&0\\ 0&1\end{bmatrix}$.
@@
(17) $\mathtt{E}^{(c)}_{\lambda=0}:=\begin{bmatrix}0&0\\ 0&1\end{bmatrix},\qquad\mathtt{E}^{(s)}_{\lambda=0}\equiv\mathtt{E}^{(s)}_{\lambda=0}(\alpha):=I_{2}-\alpha J\quad\mathrm{with}\quad J:=\begin{bmatrix}1&1\\ &1\end{bmatrix}$.
We observe an unpleasant feature of $\left.\kern-1.2ptE^{(s)}\vphantom{\big|}\right|_{\mathcal{V}_{\lambda=0}}$} : $\mathtt{E}^{(s)}_{\lambda=0}$ is no longer diagonal and its Jordan basis depends on $\alpha$. Nevertheless, a direct calculation reveals that for any $\alpha,\beta\in\mathbb{C}$ we have
@@
and so $\mathtt{E}^{(s)}_{\lambda=0}(\alpha),\mathtt{E}^{(s)}_{\lambda=0}(\beta)$ commute for any $\alpha,\beta\in\mathbb{C}$, justifying the definition
@@
$J^{i}=\begin{bmatrix}1&i\\ &1\end{bmatrix},\quad\mathtt{E}^{(c)}_{\lambda=0}J^{i}=\mathtt{E}^{(c)}_{\lambda=0},\quad\mathrm{and}\quad J^{i}\mathtt{E}^{(c)}_{\lambda=0}=\begin{bmatrix}0&i\\ 0&1\end{bmatrix}$
@@
$=\left(\mathtt{E}^{(c)}_{\lambda=0}+\sum\limits_{i=1}^{m}\beta_{i}J^{i}\mathtt{E}^{(c)}_{\lambda=0}\right)p(1)=\begin{bmatrix}0&\left(\sum\limits_{i=1}^{m}i\beta_{i}\right)\\ 0&1+\left(\sum\limits_{i=1}^{m}\beta_{i}\right)\end{bmatrix}p(1)$
$=\begin{bmatrix}0&p(1)\left(\sum\limits_{i=1}^{m}i\beta_{i}\right)\\ 0&p(1)^{2}\end{bmatrix}$.
@@
This is then true for all eigenpairs of $T$ with zero eigenvalue, thereby calculating the spectrum of $E$ restricted to the invariant subspace $\bigoplus_{\lambda=0}\mathcal{V}_{\lambda=0}$. Moreover, the eigenvectors corresponding to these eigenvalues can be explicitly evaluated. Notice also that all of the calculations above simplify notably if $D^{-1}C\mathbf{w}_{\lambda=0}=\mathbf{0}$, i.e., if $\mathbf{w}_{\lambda=0}\in\mathrm{Ker}\left(C\right)$. We summarize the results of this section below.
@@
    if $\mathtt{E}^{(s,\mathrm{full})}$ is regular. Otherwise, either $\rho^{+}$ or $\rho^{-}$ is zero and
@@
up to normalization. Pre-multiplying the eigenvectors by $V_{\lambda=0}$ gives $2(k-k_{0})$ eigenpairs of $E$ and the statement.
@@
up to the normalization. If $\mathtt{E}^{(s,\mathrm{full})}$ is singular, i.e., either $\rho^{-}$ or $\rho^{+}$ is equal to zero, then either the vector $\mathbf{e}_{1}=[1,0]^{T}$ or $\mathbf{e}_{2}=[0,1]^{T}$ lies in the kernel of $\mathbf{v}(\lambda)^{*}\mathtt{E}^{(s,\mathrm{full})}$ and hence becomes its eigenvector with eigenvalue 0. Pre-multiplying the eigenvectors by $V_{\lambda\neq 0}$ gives 2((Nn)k)2((N-n)-k) eigenpairs of $E$ and the statement.
@@
$\{0\}\cup\bigcup\limits_{\lambda\in\sigma(T)}\rho_{m}(\lambda)\subset\sigma(E),$
@@
$\{1\}\cup\bigcup\limits_{\lambda\in\sigma(T)}1-\rho_{m}(\lambda)\subset\sigma(M^{-1}L).$
Next, the question needs to be asked – can we compute the entire spectrum of $E$? To do that, we need to understand the error propagation operator $E$ also on the rest of $\mathbb{C}^{N}$ and we explore this in the next section.
@@ 3.3 Complement of the invariant subspaces
We note that by construction all $\mathbf{w}_{\lambda}$ are linearly independent and thereby $\mathcal{V}_{\lambda_{i}}\cap\mathcal{V}_{\lambda_{j}}=\delta_{ij}\mathcal{V}_{\lambda_{i}}$ and hence all of these direct sums are well defined.
@@
(21) $\mathrm{Ker}(E^{(c)})=\left\{\begin{bmatrix}\mathbf{x}_{1}\\ \mathbf{x}_{2}\end{bmatrix}\in\mathbb{C}^{N}\;|\;\mathbf{x}_{2}\in\mathbb{C}^{n}\;\mathrm{and}\;\mathbf{x}_{1}=-A^{-1}B\mathbf{x}_{2}\in\mathbb{C}^{N-n}\right\},$
@@
Let
$\mathcal{S}:=\left(\mathcal{V}^{(1)}\oplus\mathcal{V}^{(2)}\right)\cap\mathrm{Ker}(E^{(c)}),$
with $\mathcal{V}^{(1,2)}$ as in eq. (20). Then $\mathcal{S}=\mathcal{V}^{(2)}\cap\mathrm{Ker}(E^{(c)})$ and
@@
Plugging the latter into the former gives the condition $\zeta_{1}=-\lambda\zeta_{2}$ and thereby showcases that for any such $\lambda$, $\mathcal{V}_{\lambda}\cap\mathrm{Ker}(E^{(c)})$ is a one-dimensional vector space. Therefore, $\mathcal{V}^{(2)}\cap\mathrm{Ker}(E^{(c)})$ has dimension equal to the number of such $\lambda\in\sigma(T)$, i.e., to $N-n-k_{0}$, yielding the statement.
@@
The crucial observation is that the complement of $\mathcal{V}^{(1)}\oplus\mathcal{V}^{(2)}$ lies fully within $\mathrm{Ker}(E^{(c)})$. Therefore, for arbitrary smoothing parameters $\alpha_{1},\dotsc,\alpha_{m}$ the components of the error lying in the complement of $\mathcal{V}^{(1)}\oplus\mathcal{V}^{(2)}$ are zero-ed out by $E^{(c)}$ and cannot be re-introduced by the post-smoothing since $\mathcal{V}^{(1)}\oplus\mathcal{V}^{(2)}$ is an invariant subspace of $E^{(s,\mathrm{full})}$. Altogether, we have justified that we can analyze the error propagation operator $E$ purely on $\mathcal{V}^{(1)}\oplus\mathcal{V}^{(2)}$, where Theorem 3.1 gives us the full spectral information of $E$.
@@ 4 Spectral clustering
This section harvests the results of Section 3 and we show how to choose the optimal smoothing parameters $\alpha_{1},\dotsc,\alpha_{m}$. Starting with $m=1$, the nonzero eigenvalues of $E$ become
@@
From the viewpoint of choosing a suitable $\alpha$, $\rho_{1}(\alpha,\lambda)$ is a parametric polynomial whose values give the nonzero eigenvalues of $E$. Clearly we cannot nullify this polynomial by a single choice of $\alpha$ for all possible $\lambda$ to obtain a direct solver. Instead, we seek to perfectly cluster all the nonzero eigenvalues of $E$ to a single value, making it satisfy a degree 2 minimal polynomial. To do so, we make $\rho_{1}(\alpha,\lambda)$ independent of $\lambda$. The root of the derivative of $\rho_{1}(\alpha,\lambda)$ with respect to $\lambda$ is equal to $\alpha=2/3$.
@@
Equipped with Lemma 4.1, we can formulate the key result of this section.
Proposition 4.3.
Let $m\geq 1$ and take $\rho_{m}$ as in eq. (23), i.e.,
@@
where $\tau=\sqrt{\lambda}$ may take complex values. Moreover, as in Lemma 4.1 we set
@@
Using Lemma 4.1 once again for the denominator gives
@@
Due to eq. (6) we have $\sigma\left(M^{-1}L\right)=1-\sigma(E)$ and Theorem 3.1 combined with Proposition 3.3 gives
@@
The result then follows from Proposition 4.3.
@@ 5 The direct method
Denoting by $M^{-1}$ the action of Algorithm 2 as in Corollary 4.5, we have that $M^{-1}L$ is diagonalizable and has a two-point spectrum $\{1,\rho_{m}^{opt},\}$ with $\rho_{m}^{opt}\equiv 1-1/(2m+1)^{2}$. Therefore, the minimal polynomial of $M^{-1}L$ is
@@
Instead of calculating the minimal polynomial of $M^{-1}L$ explicitly, an intuitive alternative is to use a KSM to solve the system with $M^{-1}L$ – since $M^{-1}L$ is diagonalizable, both CG and GMRES will converge in at most two iterations for any right-hand side vector in exact arithmetic [^5] While the CG convergence behavior is fully described by the spectrum of the system matrix, this isn’t the case for GMRES (even in exact arithmetic; see [liesen2013krylov, Sections 5.7.2-5.7.4]). However, if the system matrix is diagonalizable with precisely $\ell$ distinct eigenvalues, then, in exact arithmetic, GMRES converges in at most $\ell$ steps., requiring at most two application of $M^{-1}$ and $L$ each. A multilevel adaptation is possible – in exact arithmetic, we can identify the two iterations of CG/GMRES for the system with $M^{-1}L$ with the inverse $(M^{-1}L)^{-1}$ at each level and we obtain a W-cycle structure, replacing the explicit formula (26) with two CG/GMRES iterations.
Since we need two iterations of a KSM using Algorithm 2 as a two-level preconditioner (inverting $M_{0}$ explicitly), a recursive, direct multilevel method takes the form of a W-cycle. Both Algorithm 1 and Algorithm 2 produce the exact solution after one iteration of their respective cycles when used as a preconditioner of a KSM, but the former is a simpler V-cycle – the W-cycle is the price to pay for the “symmetrization” of the method and regularity of the smoother. Next, we highlight two other takeaways from eq. (25), which open doors to further numerical approximations of the direct method.
In fact, “CG/GMRES to precondition CG/GMRES” is an idea that has been considered before and led to the so-called flexible GMRES method [saad1993flexible] (fGMRES) or flexible or generalized CG method [axelsson1991blackbox, notay2000flexible] (fCG, gCG). The term flexible has been used to highlight that the KSM of choice does not require the same preconditioner at each iteration; in other words the preconditioner can be non-linear.
This becomes useful for numerical modifications or approximations of (parts of) Algorithm 2, i.e., of $M^{-1}$. As a result of these approximations, two iterations of CG/GMRES no longer correspond to the exact solver and since KSM are, generally speaking, non-linear solvers we naturally obtain a non-linear preconditioner within the recursion. Therefore we need to use the flexible versions of CG/GMRES if we want to use an approximation of the multilevel version of Algorithm 2 coupled with CG/GMRES. As a result, we can still run two iterations (or more) of an optimal polynomial method, such as CG or GMRES, at each level of the recursion rather than applying a fixed polynomial as suggested by eq. (26).

### v3

@@ 3.2 Invariant subspaces – the general case
In this section, we show that the ideas of the simplified case in Section 3.1 extend to a more general setting. We assume only that $n\geq N/2$ and $T$ is diagonalizable. Zero may be an eigenvalue of $T$, and we denote its multiplicity by $k$. Eigenvectors with eigenvalue zero are denoted $\mathbf{w}_{\lambda=0}$, while all others are denoted $\mathbf{w}_{\lambda\neq 0}$.
@@
This implies that $\mathtt{E}_{\lambda\neq 0}$ is also rank-one and only one of its two eigenvalues is nonzero. Applying $\mathtt{E}_{\lambda\neq 0}$ to the vector $\mathtt{E}^{(s,\mathrm{full})}_{\lambda\neq 0}\mathbf{u}$ we get
(14) $\mathtt{E}_{\lambda\neq 0}\mathtt{E}^{(s,\mathrm{full})}_{\lambda\neq 0}\mathbf{u}$ $=\left(\mathtt{E}^{(s,\mathrm{full})}_{\lambda\neq 0}\mathbf{u}\mathbf{v}(\lambda)^{*}\mathtt{E}^{(s,\mathrm{full})}_{\lambda\neq 0}\right)\mathtt{E}^{(s,\mathrm{full})}_{\lambda\neq 0}\mathbf{u}$
@@
For each eigenpair with zero eigenvalue $(\lambda,\mathbf{w}_{\lambda=0})$ of $T$ define the subspace $\mathcal{V}_{\lambda=0}\subset\mathbb{C}^{N}$ through
(16) $V_{\lambda=0}:=\begin{bmatrix}&\mathbf{w}_{\lambda=0}\\ D^{-1}C\mathbf{w}_{\lambda=0}&\end{bmatrix}\quad\mathrm{and}\quad\mathcal{V}_{\lambda=0}=\mathrm{Im}\left(V_{\lambda=0}\right).$
@@
$E^{(s)}V_{\lambda=0}$ $=\begin{bmatrix}(1-\alpha)I&-\alpha A^{-1}B\\ -\alpha D^{-1}C&(1-\alpha)I\end{bmatrix}\begin{bmatrix}&\mathbf{w}_{\lambda=0}\\ D^{-1}C\mathbf{w}_{\lambda=0}&\end{bmatrix}$
@@
$E^{(c)}V_{\lambda=0}$ $=\begin{bmatrix}I&A^{-1}B\\ &\end{bmatrix}\begin{bmatrix}&\mathbf{w}_{\lambda=0}\\ D^{-1}C\mathbf{w}_{\lambda=0}&\end{bmatrix}=V_{\lambda=0}\begin{bmatrix}&\\ &1\end{bmatrix}$.
@@
(17) $\mathtt{E}^{(c)}_{\lambda=0}:=\begin{bmatrix}&\\ &1\end{bmatrix},\qquad\mathtt{E}^{(s)}_{\lambda=0}\equiv\mathtt{E}^{(s)}_{\lambda=0}(\alpha):=I_{2}-\alpha J\quad\mathrm{with}\quad J:=\begin{bmatrix}1&1\\ &1\end{bmatrix}$.
We observe an unpleasant feature of $\left.\kern-1.2ptE^{(s)}\vphantom{\big|}\right|_{\mathcal{V}_{\lambda=0}}$} : $\mathtt{E}^{(s)}_{\lambda=0}$ is no longer diagonal nor diagonalizable and its Jordan basis depends on $\alpha$. Nevertheless, a direct calculation reveals that for any $\alpha,\beta\in\mathbb{C}$ we have
@@
and so $\mathtt{E}^{(s)}_{\lambda=0}(\alpha)$ and $\mathtt{E}^{(s)}_{\lambda=0}(\beta)$ commute for any $\alpha,\beta\in\mathbb{C}$, justifying the definition
@@
$J^{i}=\begin{bmatrix}1&i\\ &1\end{bmatrix},\quad\mathtt{E}^{(c)}_{\lambda=0}J^{i}=\mathtt{E}^{(c)}_{\lambda=0},\quad\mathrm{and}\quad J^{i}\mathtt{E}^{(c)}_{\lambda=0}=\begin{bmatrix}&i\\ &1\end{bmatrix}$
@@
$=\left(\mathtt{E}^{(c)}_{\lambda=0}+\sum\limits_{i=1}^{m}\beta_{i}J^{i}\mathtt{E}^{(c)}_{\lambda=0}\right)p(1)=\begin{bmatrix}&\left(\sum\limits_{i=1}^{m}i\beta_{i}\right)\\ &1+\left(\sum\limits_{i=1}^{m}\beta_{i}\right)\end{bmatrix}p(1)$
$=\begin{bmatrix}&p(1)\left(\sum\limits_{i=1}^{m}i\beta_{i}\right)\\ &p(1)^{2}\end{bmatrix}$.
@@
This is true for all eigenpairs of $T$ with zero eigenvalue, thereby calculating the spectrum of $E$ restricted to the invariant subspace $\bigoplus_{\lambda=0}\mathcal{V}_{\lambda=0}$. The eigenvectors corresponding to these eigenvalues are computed in Theorem 3.1 below. Notice also that all of the calculations above simplify if $D^{-1}C\mathbf{w}_{\lambda=0}=\mathbf{0}$, i.e., if $\mathbf{w}_{\lambda=0}\in\mathrm{Ker}\left(C\right)$. We summarize the results of this section below.
@@
    provided $\mathtt{E}^{(s,\mathrm{full})}$ is regular. Otherwise, either $\rho^{+}$ or $\rho^{-}$ is zero and
@@
up to normalization. Pre-multiplying the eigenvectors by $V_{\lambda=0}$ gives the $2(k-k_{0})$ eigenpairs of $E$ and the statement.
@@
up to the normalization. If $\mathtt{E}^{(s,\mathrm{full})}$ is singular, i.e., either $\rho^{-}$ or $\rho^{+}$ is equal to zero, then either the vector $\mathbf{e}_{1}=[1,0]^{T}$ or $\mathbf{e}_{2}=[0,1]^{T}$ lies in the kernel of $\mathbf{v}(\lambda)^{*}\mathtt{E}^{(s,\mathrm{full})}$ and hence becomes its eigenvector with eigenvalue 0. Pre-multiplying the eigenvectors by $V_{\lambda\neq 0}$ gives the 2((Nn)k)2((N-n)-k) eigenpairs of $E$ and the statement.
@@
$\left\{\{0\}\cup\bigcup\limits_{\lambda\in\sigma(T)}\rho_{m}(\lambda)\right\}\subset\sigma(E),$
@@
$\left\{\{1\}\cup\bigcup\limits_{\lambda\in\sigma(T)}1-\rho_{m}(\lambda)\right\}\subset\sigma(M^{-1}L).$
To compute the entire spectrum of $E$, we need to understand the error propagation operator $E$ on the rest of $\mathbb{C}^{N}$; we explore this in Section 4. The next section is dedicated to complete the description of the eigenspaces.
@@ 3.3 Complement of the invariant subspaces
We note that, thanks to our assumption that $T$ is diagonalizable, the vectors $\mathbf{w}_{\lambda}$ are linearly independent, $\mathcal{V}_{\lambda_{i}}\cap\mathcal{V}_{\lambda_{j}}=\delta_{ij}\mathcal{V}_{\lambda_{i}}$, and therefore all of these direct sums are well defined.
@@
(21) $\mathrm{Ker}(E^{(c)})=\left\{\begin{bmatrix}\mathbf{x}_{1}\\ \mathbf{x}_{2}\end{bmatrix}\in\mathbb{C}^{N}\;|\;\mathbf{x}_{2}\in\mathbb{C}^{n}\;\mathrm{and}\;\mathbf{x}_{1}=-A^{-1}B\mathbf{x}_{2}\in\mathbb{C}^{N-n}\right\}$
@@
Let $\mathcal{S}:=\left(\mathcal{V}^{(1)}\oplus\mathcal{V}^{(2)}\right)\cap\mathrm{Ker}(E^{(c)})$, with $\mathcal{V}^{(1,2)}$ as in eq. (20). Then $\mathcal{S}=\mathcal{V}^{(2)}\cap\mathrm{Ker}(E^{(c)})$ and
@@
Plugging the latter into the former gives the condition $\zeta_{1}=-\lambda\zeta_{2}$, thereby showcasing that for any such $\lambda$, $\mathcal{V}_{\lambda}\cap\mathrm{Ker}(E^{(c)})$ is a one-dimensional vector space. Therefore, $\mathcal{V}^{(2)}\cap\mathrm{Ker}(E^{(c)})$ has dimension equal to the number of such $\lambda\in\sigma(T)$, i.e., to $N-n-k_{0}$, yielding the statement.
@@
The crucial observation is that the complement of $\mathcal{V}^{(1)}\oplus\mathcal{V}^{(2)}$ lies fully within $\mathrm{Ker}(E^{(c)})$. Therefore, for arbitrary smoothing parameters $\alpha_{1},\dotsc,\alpha_{m}$ the components of the error lying in the complement of $\mathcal{V}^{(1)}\oplus\mathcal{V}^{(2)}$ are zero-ed out by $E^{(c)}$ and cannot be re-introduced by the post-smoothing since $\mathcal{V}^{(1)}\oplus\mathcal{V}^{(2)}$ is an invariant subspace of $E^{(s,\mathrm{full})}$. Altogether, we have justified that we can analyze the error propagation operator $E$ purely on $\mathcal{V}^{(1)}\oplus\mathcal{V}^{(2)}$, where Theorem 3.1 gives us the full spectral information of $E$. The next section will provide the exact choice of $\alpha_{1},\dotsc,\alpha_{m}$ such that the spectrum of $E$ collapses to a single nonzero eigenvalue for any choice of $m$.
@@ 4 Spectral clustering
In this section we harvest the results of Section 3 and compute the optimal smoothing parameters $\alpha_{1},\dotsc,\alpha_{m}$. Starting with $m=1$, the nonzero eigenvalues of $E$ become
@@
From the viewpoint of choosing a suitable $\alpha$, $\rho_{1}(\alpha,\lambda)$ is a parametric polynomial whose values give the nonzero eigenvalues of $E$. Clearly we cannot nullify this polynomial by a single choice of $\alpha$ for all possible $\lambda$ (and thereby obtain a direct solver). Instead, we seek to perfectly cluster all the nonzero eigenvalues of $E$ to a single value, making it satisfy a degree 2 minimal polynomial. To do so, we make $\rho_{1}(\alpha,\lambda)$ independent of $\lambda$: we compute the root of the derivative of $\rho_{1}(\alpha,\lambda)$ with respect to $\lambda$, obtaining $\alpha=2/3$.
@@
Equipped with Prop. 4.1, we can formulate the key result of this section.
Lemma 4.3.
Let $m\geq 1$, take $\rho_{m}$ as in eq. (23), i.e.,
@@
where $\tau=\sqrt{\lambda}$ may take complex values, and let, as in Prop. 4.1,
@@
Using Prop. 4.1 once again for the denominator gives
@@
Due to eq. (6) we have $\sigma\left(M^{-1}L\right)=1-\sigma(E)$. Theorem 3.1 combined with Proposition 3.3 gives
@@
The result then follows from Lemma 4.3.
@@ 5 The direct method
Denoting by $M^{-1}$ the action of Algorithm 2 as in Theorem 4.5, we have that $M^{-1}L$ is diagonalizable and has a two-point spectrum $\{1,\rho_{m}^{opt},\}$ with $\rho_{m}^{opt}\equiv 1-1/(2m+1)^{2}$. Therefore, the minimal polynomial of $M^{-1}L$ is
@@
Instead of calculating the minimal polynomial of $M^{-1}L$ explicitly, an intuitive alternative is to use a KSM to solve the system with $M^{-1}L$ – since $M^{-1}L$ is diagonalizable, both CG and GMRES will converge in at most two iterations for any right-hand side vector in exact arithmetic [^4] While the CG convergence behavior is fully described by the spectrum of the system matrix, this isn’t the case for GMRES (even in exact arithmetic; see [liesen2013krylov, Sections 5.7.2-5.7.4]). However, if the system matrix is diagonalizable with precisely $\ell$ distinct eigenvalues, then, in exact arithmetic, GMRES converges in at most $\ell$ steps., requiring at most two applications of $M^{-1}$ and $L$ each. A multilevel adaptation is possible – in exact arithmetic, we can identify the two iterations of CG/GMRES (with $M^{-1}L$ as the system matrix) with the inverse $(M^{-1}L)^{-1}$ at each level, replacing the explicit formula (26) with two CG/GMRES iterations at each level.
Since we need two iterations of a KSM using Algorithm 2 as a two-level preconditioner (inverting $M_{0}$ explicitly), a recursive, direct multilevel method takes the form of a K-cycle with two recursive coarse corrections, thus also a W-cycle. Both Algorithm 1 and Algorithm 2 produce the exact solution after one iteration of their respective cycles when used as a preconditioner of a KSM, but the former is a simpler V-cycle – the K/W-cycle is the price to pay for the “symmetrization” of the method and regularity of the smoother. Next, we highlight two other takeaways from eq. (25), which open doors to further numerical approximations of the direct method.
Using CG/GMRES to precondition CG/GMRES has been considered before and led to the flexible GMRES method [saad1993flexible] (fGMRES) and flexible or generalized CG method [axelsson1991blackbox, notay2000flexible] (fCG, gCG). The term flexible has been used to highlight that the KSM of choice does not require the same preconditioner at each iteration; the preconditioner can be nonlinear.
This becomes useful for numerical modifications or approximations of (parts of) Algorithm 2, i.e., of $M^{-1}$. As a result of these approximations, two iterations of CG/GMRES no longer correspond to the exact solver. Since KSMs are generally nonlinear solvers, we naturally obtain a nonlinear preconditioner within the recursion. Therefore, we need to use the flexible versions of CG/GMRES if we want to use an approximation of the multilevel version of Algorithm 2 coupled with CG/GMRES. As a result, we can still run two iterations (or more) of an optimal polynomial method, such as CG or GMRES, at each level of the recursion rather than applying a fixed polynomial as suggested by eq. (26).